OpenStudy (anonymous):

Integral problem: Let \(f_n(x) =\begin{cases} \dfrac{1}{x}~~~\text{if}~~~ x \in [1/n, 1]\\ 0, ~~~\text{otherwise} \end{cases}\), for every \(n\) in \(\mathbb{N}\). Is \(f_n\) Lebesgue integrable?

OpenStudy (anonymous):

I'm pretty sure it is not, since at some point i'd have \(n \rightarrow \infty\) \(\Rightarrow [1/n, 1] \rightarrow [0, 1]\) and this is definately not Riemann or Lebesgue integrable (if you look at the improper Riemann integral) Even if only had \((0,1]\) I'd still have the same problem. (or am I wrong)?

OpenStudy (thomas5267):

I don't know anything about Lebesgue integral but I think it might even be Riemann integrable for every \(n\in\mathbb{N}\setminus \{0\}\). For every \(n\), the \(f_n\) has finite support and the integral is finite as well, and \(\infty\notin \mathbb{N}\).

OpenStudy (anonymous):

Ooh yeah, so infinity isn't in the Naturals. It seems like that completely went over my head, then everything makes a lot more sense :)

OpenStudy (anonymous):

Thanks a lot for that.